OpenStudy (anonymous):

Find the value of the sec 19° using your calculator. 1.) 0.946 2.) 1.011 3.) 0.989 4.) 1.058

Nnesha (nnesha):

what's the definition of sec theta ??? like csc =1/sin theta

Nnesha (nnesha):

nvm \[\sin \theta =\frac{ 1 }{ \csc \theta }~~~\cos \theta =\frac{ 1 }{ \sec \theta }~~~\tan \theta =\frac{ 1 }{\cot \theta }\] sin,cos ,tan and csc ,sec ,cot are reciprocal if sin = 1/csc then csc = 1/sin same way if cos =1/sec then sec =1/cos now plugin to the calculator ....